# Performance evaluation of compact plastic scintillating fiber modules for muon tomography applications

### v1

@@ 4.2 Signal clustering and track reconstruction
$x_{\mathrm{cog}}=\frac{\sum_{i}N_{i}x_{i}}{\sum_{i}N_{i}},$ (4.2)
where $x_{i}$ is the center position of the $i$th SiPM in the cluster, and $N_{i}$ is its corresponding signal.
To suppress background clusters or clusters contaminated by background channels, only those with a cluster size $cls\leq 3$ in one layer are retained for track reconstruction. Particle tracks are reconstructed separately in the $xz$ and $yz$ planes. All possible combinations of clusters across different layers are scanned and tested with a straight-line fit by the least square method, and the set of clusters yielding the smallest fitting $\chi^{2}$ is selected as the track candidate for that event. Fig. 6 illustrates an event display of reconstructed tracks in the $xz$ and $yz$ projection planes after alignment as explained in Sec. 4.3.
@@ 4.3 Module alignment
$R^{2}=\sum_{N}\sum_{i}\left[\left(y_{i}^{c}-y_{i}^{f}\right)^{2}+\left(x_{i}^{c}-x_{i}^{f}\right)^{2}\right],$ (4.3)
@@
During the alignment procedure, only the cluster with the maximum signal on each layer is used in track fitting. Tracks are required to satisfy the following criteria: a cluster size of $cls\leq 3$ on all four layers, and a fit quality of $\chi^{2}<100$. The tracks reconstructed with cluster of $cls=1$ on all four long or short modules are dimissed to avoid biased residuals as explained in Sec. 5.2. The minimization of Eq. 4.3 is performed iteratively with the Gauss–Newton method. For each test position of SciFi modules, at least 1000 selected tracks are used for the alignment. Covergence to level of 0.1 mm of translation parameters is realized typically after 3 iterations. Fig. 7 shows an example of the residual distributions with tracks of $\chi^{2}<10$ of all SciFi modules before and after alignment using muon test data, clearly demonstrating that the mean of the residuals shifts to zero after alignment.
5 Results
The detection efficiency and spatial resolution are key performance parameters of the SciFi modules, as they determine both the imaging time and the accuracy of muon tomography. These parameters are evaluated using muon events recorded at five different distances between the center of the short module and the readout end of the long SciFi modules along the $x$ axis. The different measurement positions were achieved by sliding the short modules within the experimental setup. For each position, data are collected over periods longer than five days in muon-test mode.
@@ 5.1 Detection efficiency
To evaluate the detection efficiency of a long SciFi module under test (DUT) at a given measurement position, good events are selected according to the following criteria:
  • The short modules, excluding the one paired with the DUT, are required to reconstruct $xz$-plane tracks with $\chi^{2}<2$ and cluster sizes of $cls\leq 3$ in each participating layer. This ensures that the selected muons traversed the intended measurement position on the long modules.
  • The long modules, excluding the DUT, are required to reconstruct $yz$-plane tracks with $\chi^{2}<2$ and cluster sizes of $cls\leq 3$ in each participating layer. In addition, the extrapolated track position on the DUT is required to lie within its sensitive area.
@@
With these criteria, about 1000 good events per day are obtained for each DUT. Fig. 8 shows, as an example, the cluster size and cluster signal distributions for the DUT in layer 2 at a measurement position of 52.5 cm. Most clusters have $cls\leq 2$, and the most probable value of $N_{\mathrm{pe}}$ is about 36 p.e., consistent with the module design expectation. Clusters with larger sizes are typically associated with higher signals, which can originate from muon hit combined with secondary particles produced in muon interactions, such as $\delta$ rays, or with ambient radiation.
For each good event, the DUT is considered to have detected the muon if a cluster with $cls\leq 3$ is present on the DUT layer and the residual between the cluster position and the extrapolated track position in the $yz$ plane is less than 5 mm. Events with $cls=3$ are retained to maximize statistics while only marginally affecting the position resolution. The residual threshold is particularly relevant for the outermost modules, where it is influenced by the geometrical layout. The detection efficiency is then defined as
$\epsilon=\frac{N_{\mathrm{det}}}{N_{\mathrm{good}}}\;,$ (5.1)
where $N_{\mathrm{det}}$ is the number of events with a valid cluster on the DUT and $N_{\mathrm{good}}$ is the total number of good events. Fig. 9 shows the measured detection efficiencies of the four long SciFi modules at different measurement positions. The efficiencies are uniform along the module length and consistently exceed 97%.
@@ 5.2 Spatial resolution
The spatial resolution of a DUT is evaluated from the residual distribution between the detected cluster position on the DUT and the extrapolated track reconstructed with the other layers. The measured residual variance, $\sigma_{r}$, includes contributions from both the intrinsic spatial resolution of the DUT, $\sigma_{d}$, and the resolution of the extrapolated track, $\sigma_{f}$:
$\sigma_{r}^{2}=\sigma_{d}^{2}+\sigma_{f}^{2}\;.$ (5.2)
Since tracks are reconstructed with a straight-line fit, the extrapolated hit position is a linear combination of the reference layers. Accordingly, the resolution of the extrapolated position at module $m$ can be expressed as [22]:
$\sigma_{f,m}^{2}$ $=\sum_{j\neq m}w_{mj}\sigma_{d,j}^{2},$ (5.3)
@@
where $N$ is the total number of long modules and $z_{j}$ denotes the vertical position of module $j$ along the $z$ axis. The spatial resolution of each module was then obtained by solving the set of linear equations for all four modules.
@@
Because the SiPM pitch is larger than the fiber diameter, about 68% of muon clusters correspond to $cls=1$, as shown in Fig. 8. In these cases, the reconstructed cluster positions by the CoG method are fixed at the SiPM channel centers. As a result, roughly 30% of tracks are reconstructed exclusively from $cls=1$ clusters in only three reference layers. Such tracks with hits on three SiPM centers lead to systematic reconstruction biases imposed by the fixed detector layout of SiPMs as illustrated in Fig. 6. Furthermore, since only DUT clusters with residuals smaller than 5 mm are accepted, the residual distribution of such biased tracks shows discrete spikes for $cls=1$ clusters, as seen in Fig. 10(b). Similar patterns are also observed for other layers.
After excluding tracks reconstructed solely from $cls=1$ clusters in the reference layers, the combined residual distribution of DUT clusters with $cls=1$ and $cls>1$ is well described by a Gaussian fit, from which $\sigma_{r}$ is extracted (Fig. 10(c)). Although the intrinsic resolution is expected to be better for $cls>1$ clusters than for $cls=1$, in analogy with the performance of segmented pixel detectors in Ref. [23], here we report only the overall spatial resolution constrained by the limited number of experimental layers and muon events. Fig. 11 shows the extracted spatial resolutions of all four SciFi modules at different measurement positions. The average resolution across all layers and positions is about $0.56\pm 0.04$ mm. The modules exhibit an approximately uniform response along their length.
@@ 6 Conclusion
In this work, we designed and constructed four compact 100 cm long SciFi modules, each comprising two staggered layers of 1 mm diameter plastic scintillating fibers read out at one end by a one-dimensional SiPM array with a 2 mm pitch. The modules were characterized using cosmic-ray muons triggered by 10 cm short SciFi modules positioned at different distances along their length, with data acquired by Citiroc 1A–based electronics. After dedicated channel calibration, signal clustering, track reconstruction, and alignment, the modules achieved detection efficiencies above 97% and an average spatial resolution of about 0.56 mm. Both parameters were found to be uniform across the full module length.
Compared with previously developed scintillating fiber detectors for muon tomography, the present design is more compact and avoids the laborious fiber-grouping process, while providing sub-millimeter spatial resolution superior to that of bulk scintillator detectors. These features demonstrate the potential of the proposed modules for building compact, high-resolution muon tomography systems. Future work focusing on SiPM readout multiplexing strategies is under development to reduce the number of electronics channels, thereby advancing toward scalable and cost-effective large-scale applications.
Acknowledgments

### v2

@@ 4.2 Signal clustering and track reconstruction
$x_{\mathrm{cog}}=\frac{\sum_{i}N_{i}x_{i}}{\sum_{i}N_{i}},$ (2)
where $x_{i}$ is the center position of the $i$th SiPM in the cluster, and $N_{i}$ is its corresponding signal size.
Typically, muon events may result in signals from one or two channels in the SciFi module. Crosstalk between adjacent fibers due to imperfect glue coverage may also induce signals from $cls=3$, which was observed during quality inspections of the modules. To suppress background clusters or clusters contaminated by background channels, only those with a cluster size $cls\leq 3$ in a given layer are retained for track reconstruction. Particle tracks are reconstructed separately in the $xz$ and $yz$ planes. All possible combinations of clusters across different layers are scanned and tested with a straight-line fit using the least squares method, and the set of clusters yielding the smallest fitting $\chi^{2}$ is selected as the track candidate for that event. Fig. 7 illustrates an event display of reconstructed tracks in the $xz$ and $yz$ projection planes after alignment, as explained in Sec. 4.3. Fig. 8 shows the two-dimensional angular distribution of reconstructed tracks after alignment to illustrate the field of view of the entire setup.
@@
Figure 8: The angular map of reconstructed tracks after alignment. Here the $yz$ tracks are reconstructed with layers 1,3,4, and the $xz$ tracks are reconstructed with layers 5,7,8.
@@ 4.3 Module alignment
$R^{2}=\sum_{N}\sum_{i}\left[\left(y_{i}^{c}-y_{i}^{f}\right)^{2}+\left(x_{i}^{c}-x_{i}^{f}\right)^{2}\right],$ (3)
@@
During the alignment procedure, only the cluster with the maximum signal on each layer is used in track fitting. Tracks are required to satisfy the following criteria: a cluster size of $cls\leq 3$ on all four layers, and a fit quality of $\chi^{2}<100$. The tracks reconstructed with cluster of $cls=1$ on all four long or short modules are dismissed to avoid biased residuals as explained in Sec. 5.2. The minimization of Eq. 3 is performed iteratively with the Gauss–Newton method [Blobel2007]. For each test position of SciFi modules, at least 1000 selected tracks are used for the alignment. Convergence to level of 0.1 mm of translation parameters is realized typically after 3 iterations. Fig. 9 shows an example of the residual distributions with tracks of $\chi^{2}<10$ of all SciFi modules before and after alignment using muon test data, clearly demonstrating that the mean of the residuals shifts to zero after alignment.
5 Module performance
The detection efficiency and spatial resolution are key performance parameters of the SciFi modules, as they determine both the imaging time and the reconstruction accuracy in muon tomography. In this section, we take a single long SciFi module as the Device Under Test (DUT) and evaluate these parameters for each DUT individually. Measurements are performed at five positions along the $x$ axis, defined as the distance between the center of the short module and the readout end of the long module. These positions are realized by sliding the short modules within the setup. For each position, data are collected over periods longer than five days in muon-test mode.
@@ 5.1 Detection efficiency
To determine the detection efficiency of the DUT, we define $\rm{N_{good}}$ as the number of muon tracks expected to pass through the fiducial volume of the DUT, and $\rm{N_{det}}$ as the number of these tracks that produce a valid cluster on the DUT. Fig. 10 illustrates an example that the detection efficiency calculation for layer 2 as the DUT, where the short modules except for the one paired with the DUT are employed to constrain the muon impact position, while the other long modules are used to reconstruct the reference tracks for efficiency determination.
Figure 10: Illustration of the detection efficiency calculation for a single long SciFi module (shown in red), with muon tracks selected using the short modules (shown in yellow) and the other long modules (shown in blue). In this example, layer 2 in red is treated as the DUT.
To be explicit, the selection of good muon events follows these criteria:
  • 1.
    For the short modules (excluding the one paired with the DUT), $xz$-plane tracks are reconstructed with a goodness-of-fit requirement of $\chi^{2}<2$. Clusters used in the reconstruction must have a signal of $N_{\mathrm{pe}}>8$ p.e. and a cluster size of $cls\leq 3$ in each participating layer. These conditions ensure that the selected muons traverse the intended measurement region on the long modules.
  • 2.
    For the long modules (excluding the DUT), $yz$-plane tracks are reconstructed with $\chi^{2}<2$, and only clusters with $N_{\mathrm{pe}}>8$ p.e. and $cls\leq 3$ are accepted in each layer. Furthermore, the extrapolated track position on the DUT is required to fall within its active sensitive area.
@@
With these criteria, about 1000 good events per day are obtained for each DUT. Fig. 11 shows, as an example, the cluster size and cluster signal distributions for the DUT in layer 2 at a measurement position of 52.5 cm. Most clusters have $cls\leq 2$, and the most probable value of $N_{\mathrm{pe}}$ is approximately 36 p.e., consistent with the module design expectation. Clusters with larger sizes are typically associated with higher signals, which can originate from muon hit combined with secondary particles produced in muon interactions, such as $\delta$ rays, or with ambient radiation.
For each good event, the DUT is considered to have successfully detected the muon if a cluster with $cls\leq 3$ is present on the DUT layer and the residual between the cluster position and the extrapolated track position in the $yz$ plane is less than 5 mm. Events with $cls=3$ are retained to maximize statistics while only marginally affects the position resolution. The residual threshold is particularly relevant for the outermost modules due to geometrical constraints. The detection efficiency is then defined as
$\epsilon=\frac{N_{\mathrm{DUT}}}{N_{\mathrm{good}}}\;,$ (5)
where $N_{\mathrm{DUT}}$ is the number of events with a valid cluster on the DUT and $N_{\mathrm{good}}$ is the total number of good events.
Fig. 12 shows the measured detection efficiencies of the four long SciFi modules at different measurement positions. The efficiencies are uniform along the module length and consistently exceed 97%. To further verify that this measurement reflects the absolute detection efficiency, we varied the cluster signal threshold used to select good events (8, 15, 20, and 25 p.e.). The resulting detection efficiencies remain stable across these thresholds, confirming that the efficiency evaluation is robust.
@@ 5.2 Spatial resolution
The spatial resolution of the long SciFi modules is primarily determined by the fiber diameter, the SiPM pitch, and the threshold settings used in the cluster-finding procedure. To qualitatively assess their influence, a preliminary Monte Carlo simulation of a single SciFi module was performed using the Geant4 toolkit [GEANT4:2002zbu]. Cosmic-ray muons were generated within the same zenith-angle range as in the experimental setup, following the modified Gaisser formula reported in Ref. [guan2015parametrization]. Detailed detector effects including SiPM dark noise, optical crosstalk between adjacent fibers, and the air gap between the fiber end faces and the SiPM array were not included.
The same cluster-finding algorithm described in Sec. 4.2 was applied, and the simulated cluster signal distributions were tuned to reproduce the most probable values observed in Fig. 11 by adjusting the light yield of the scintillating fibers. For each event, the residual between the reconstructed cluster position and the true hit position was evaluated under four different cluster-threshold configurations, as shown in Fig. 13. Both the full width at half maximum (FWHM) of the residual distribution and the fractions of events with $cls=1$ and $cls=2$ exhibit dependence on the neighboring threshold. As the neighboring threshold is lowered, the FWHM and the fraction of $cls=1$ events decrease, while the proportion of $cls=2$ events increases. The residual plateau width for $cls=1$ can be smaller than the SiPM pitch (2 mm), because muons traversing near the boundary between two SiPM channels tend to produce signals in two adjacent channels and are therefore categorized as $cls=2$.
[…]
Figure 13: Residual distributions of simulated events with $cls=1$ and $cls=2$ by using different threshold values during cluster finding.
In the experimental analysis, we use the cluster thresholds defined in Sec. 4.2 and the good muon events selected in Sec. 5.1. The spatial resolution of a DUT is evaluated from the residual distribution between the detected cluster position on the DUT and the extrapolated track reconstructed with the other layers. The measured residual variance, $\sigma_{r}^{2}$, includes contributions from both the intrinsic spatial resolution of the DUT, $\sigma_{d}$, and the resolution of the extrapolated track, $\sigma_{f}$:
$\sigma_{r}^{2}=\sigma_{d}^{2}+\sigma_{f}^{2}\;.$ (6)
Since tracks are reconstructed with a straight-line fit, the extrapolated hit position is a linear combination of the reference layers. Accordingly, the resolution of the extrapolated position at module $m$ can be expressed as [Wang:2022TNS]:
$\sigma_{f,m}^{2}$ $=\sum_{j\neq m}w_{mj}^{2}\sigma_{d,j}^{2},$ (7)
@@
where $N$ is the total number of long modules, $z_{j}$ denotes the vertical position of module $j$ along the $z$ axis and $w_{mj}$ refers to the weight of each layer uncertainty extracted from the error progation of straight line fit formulae. The spatial resolution of each module was then obtained by solving the set of linear equations for all four modules.
@@
Because the SiPM pitch is larger than the fiber diameter, about 68% of muon clusters correspond to $cls=1$, as shown in Fig. 11. In these cases, the reconstructed cluster positions by the CoG method are fixed at the SiPM channel centers. As a result, roughly 30% of tracks are reconstructed exclusively from $cls=1$ clusters in only three reference layers. Such tracks with hits on three SiPM centers lead to systematic reconstruction biases imposed by the fixed detector layout of the SiPMs as illustrated in Fig. 7. Furthermore, since only DUT clusters with residuals smaller than 5 mm are accepted, the residual distribution of such biased tracks shows discrete spikes for $cls=1$ clusters, as seen in Fig. 14(b). Similar patterns are also observed for other layers.
@@
Figure 16: Spatial resolution of events with $cls=1$ and $cls>1$ on layer 3.
After excluding tracks reconstructed solely from $cls=1$ clusters in the reference layers, the combined residual distribution of DUT clusters with $cls=1$ and $cls>1$ is well described by a Gaussian fit, from which $\sigma_{r}$ is extracted (Fig. 14(c)). By solving Eq. 7, the intrinsic spatial resolutions including contributions from both $cls=1$ and $cls>1$ events were obtained for each module. Figure 15 summarizes the spatial resolutions of all four SciFi modules at different measurement positions. The average resolution over all modules and positions is approximately $0.56\pm 0.04$ mm and the modules exhibit an approximately uniform response along their length. As an illustration, Fig. 16 shows the intrinsic spatial resolutions for $cls=1$ and $cls>1$ events in layer 3, obtained using the same procedure with the combined position resolutions of the other layers fixed to the values extracted previously.
6 Imaging experiment
Cosmic-ray muons passing through materials of different atomic numbers experience distinct multiple scattering angles, which can be approximated by the formula:
$\theta_{0}=\frac{13.6~\mathrm{MeV}}{\beta p}z\sqrt{\frac{x}{X_{0}}}\left[1+0.038\ln\left(\frac{x}{X_{0}}\right)\right],$ (8)
where $\theta_{0}$ is the RMS of the projected multiple scattering angle, $p$ is the particle momentum in MeV/$c$, $\beta$ is the particle velocity relative to the speed of light, $z$ is the muon charge number, $x$ is the material thickness, and $X_{0}$ is the radiation length of the material. By measuring the deflection angles of muons before and after traversing a target, materials with different atomic numbers can thus be discriminated based on muon scattering.
In the imaging setup shown in Fig. 17, the upper detector consists of layers 1, 2, 5, and 6, while the lower detector consists of layers 3, 4, 7, and 8. The separation $L$ between the two long modules in either the upper or lower detector is 10 cm. The upper and lower detectors are used to record muon trajectories before and after the target, respectively. The sensitive acceptance corresponding to a fiducial volume of approximately $\rm 5.5~cm\times 5.5~cm\times 30~cm$ is relatively small. Thus, the following imaging test serves mainly as a proof-of-principle demonstration of the capabilities of the developed modules for muon tomography.
Figure 17: Muon scattering tomography setup with a $\rm{3~cm\times 3~cm\times 6~cm}$ lead block placed between the upper and lower detectors.
As evaluated in Sec. 5, the mean position resolution $\sigma_{d}$ of a long SciFi module is approximately 0.56 mm. The corresponding angular resolution $\sigma_{\theta}$ of either the upper or lower detector is estimated as $\sigma_{\theta}=\sqrt{2}\sigma_{d}/L\approx 8~\mathrm{mrad}$. For a 4 GeV muon traversing a 6 cm thick lead block, the expected RMS scattering angle $\theta_{0}$ is about 12 mrad. Therefore, the imaging setup is expected to distinguish a $\rm{3~cm\times 3~cm\times 6~cm}$ lead block placed between the upper and lower detectors in air.
After about ten days of data acquisition in muon-test mode, roughly 4000 events containing valid tracks in both the upper and lower detectors were recorded. Point of Closest Approach (PoCA)  [schultz2003cosmic] points were reconstructed for events with deviation angles greater than 6 mrad. A $k$-Nearest Neighbors (k-NN) algorithm was subsequently applied to count the number of neighboring points within a defined distance around each PoCA point, and those with a small number of neighbors were removed as noise outliers.
The reconstructed projections are presented in Fig. 18, where color of each point represents the corresponding event’s deflection angle. Most PoCA points with large deviation angles cluster within the true lead block region. The $\rm{XY}$ projections clearly reveal the block region, the ambiguity of PoCA points along the $z$ axis in both the $\rm{XZ}$ and $\rm{YZ}$ projections mainly results from the limited geometrical acceptance of the imaging setup, where muons undergoing large deflections are more likely to fall outside the lower detector acceptance [luo2022development]. Overall, the results demonstrate the feasibility of the developed SciFi detector modules for compact muon tomography imaging applications.
[…]
Figure 18: Reconstructed PoCA scattering maps projected onto the $\rm{XY}$ (a), $\rm{XZ}$ (b), and $\rm{YZ}$ (c) planes. The color of each point represents the corresponding event’s deviation angle.
@@ 7 Conclusion
In this work, we designed and constructed four compact 100 cm long SciFi modules, each comprising two staggered layers of 1 mm diameter plastic scintillating fibers read out at one end by a one-dimensional SiPM array with a 2 mm pitch. The modules were characterized using cosmic-ray muons triggered by 10 cm short SciFi modules positioned at different distances along their length, with data acquired by Citiroc 1A–based electronics. After dedicated channel calibration, signal clustering, track reconstruction, and alignment, the modules achieved detection efficiencies above 97% and an average spatial resolution of about 0.56 mm. Both parameters were found to be uniform across the full module length. Building upon these characterization results, a proof-of-principle imaging experiment was conducted using the SciFi modules arranged in a muon scattering tomography configuration. The reconstructed image obtained via the PoCA method revealed the profile of the target lead block, demonstrating the system’s capability for muon-based imaging.
Compared with previously developed scintillating fiber detectors for muon tomography, the present design is more compact and avoids the laborious fiber-grouping process, while providing sub-millimeter spatial resolution superior to that of bulk scintillator detectors. These features demonstrate the potential of the proposed modules for building compact, high-resolution muon tomography systems. Future work focusing on the muon tomography prototype with multiplexed SiPM readout is under development toward compact and cost-effective large-scale applications.
Acknowledgements